OpenStudy (anonymous):

∠A and ∠B are complementary angles of right triangle ABC, cos A = 0.83, and cos B = 0.55. What is sin A + sin B?

OpenStudy (anonymous):

\[\sin A=\sqrt{1-\cos ^2 A},\sin B=\sqrt{1-\cos ^2B}\]

OpenStudy (anonymous):

hey!!!